# references.bib

### v2

@@ 4.   Block Coordinate Network Simplex Methods with Basis Variable Succession
  • $\mathbf{x}^{*}$:the optimal solution to the full OT problem (2),
@@ 4.1 Basis Variables Succession
The following theorem identifies the precise conditions under which a restricted NS step on a block $\mathcal{H}_{k}$ yields a globally optimal solution. It also shows when global optimality cannot be achieved.
@@
Apply the network–simplex method to the subproblem restricted to $\mathcal{H}_{k}$, and set
@@
    If $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$ and $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the iterate $\mathbf{x}^{k+1}$ is globally optimal.
@@
    If $\mathcal{B}_{k}\not\subseteq\mathcal{H}_{k}$, then—even when $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$—the iterate $\mathbf{x}^{k+1}$ cannot be globally optimal.
@@
(i) Because $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$, the NS subproblem starts with basis $\mathcal{B}_{k}$. Pivoting within $\mathcal{H}_{k}$ yields a basis ^$\widehat{\mathcal{B}}\subseteq\mathcal{H}_{k}$ satisfying
[…]
Since $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the subproblem contains an optimal basis of the full problem; NS therefore stops only when no negative reduced costs remain in $\mathcal{H}_{k}$. As dual variables depend solely on the basis,
^$\widehat{\mathcal{B}}\cap\mathcal{H}_{k}=\mathcal{B}^{\ast},$
so condition (5) extends to all indices $i=1,\dots,N$:
[…]
By complementary slackness, $\mathbf{x}^{k+1}$ is globally optimal.
(ii) Assume $j\in\mathcal{B}_{k}\setminus\mathcal{H}_{k}$ with $\mathbf{x}^{k}_{j}>0$. Because $j\notin\mathcal{H}_{k}$, the subproblem cannot pivot $j$ or alter its value, hence
$\mathbf{x}^{k+1}_{j}=\mathbf{x}^{k}_{j}>0.$
In any optimal OT solution all non–basic variables are zero. Since $j\notin\mathcal{B}^{\ast}$ (because $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$), optimality would require $x^{\ast}_{j}=0$, contradicting $\mathbf{x}^{k+1}_{j}>0$. Therefore $\mathbf{x}^{k+1}$ cannot be optimal. ∎
The above analysis motivates a practical strategy, which we call basis variable succession: at each iteration, we selecte the block $\mathcal{H}_{k}$ includes the basis variables $\mathcal{B}_{k}$ obtained from the previous subproblem. This design contrasts with existing BCD approaches [xie2023randomized] that select blocks randomly.
we can solve the subproblem in iteration $k$ with complexity $O(L_{k}|\mathcal{H}_{k}|)$, which is significantly lower than the $O(LN)$ complexity of solving the full problem directly, where $L_{k}$ and $L$ denote the number of NS iterations for the subproblem and the full problem, respectively.
This strategy offers two practical advantages:
  1. 1.
    The subproblem always includes all constraints and at least $M-1$ variables, ensuring feasibility.
  2. 2.
    Reusing the previous basis variables as the initialization eliminates the need for recomputing a starting basis in the NS method.
@@ 4.2 A Prototype of Block Coordinate Descent Network Simplex Methods and its Finite Termination
Following the strategy of §4.1, choose a block
@@
Under condition (6) each subproblem produces a strictly improving search direction, ensuring monotone decrease of the objective value. The next theorem establishes finite termination and global optimality of the resulting sequence $\{\mathbf{x}^{k}\}$.
@@
The proposed prototype of BCDNS method satisfies the following properties:
  1. 1.
    Monotonicity and Convergence: The objective function decreases monotonically at each iteration, and the method converges to the global optimal solution.
  2. 2.
    Finite Termination: The method terminates within a finite number of iterations.
Proof of Part 1.
Monotonicity: For any iteration $k$, based on the selection rule for $\mathcal{I}_{k}$, there exists at least one $l\in\mathcal{I}_{k}$ such that the reduced cost of $x_{l}$ is negative. By the properties of the reduced cost, adding $x_{l}$ to the basis variables decreases the total cost, i.e.,
$q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\{l\})<q(\mathbf{x}^{k};\mathcal{B}_{k}).$
where $q(\mathbf{x}^{k};\mathcal{I}_{k})$ is defined by (4) as the optimal value of the subproblem associated with the index set $\mathcal{I}_{k}$. Since $l\in\mathcal{I}_{k}$, it follows that:
$q(\mathbf{x}^{k};\mathcal{H}_{k})=q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\mathcal{I}_{k})\leq q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\{l\})<q(\mathbf{x}^{k};\mathcal{B}_{k}).$
Furthermore, since $\mathcal{J}_{k}$ is the optimal basis variables of the previous subproblem, we have:
$q(\mathbf{x}^{k};\mathcal{B}_{k})=q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$
Combining the above, it follows that:
$q(\mathbf{x}^{k};\mathcal{H}_{k})<q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$
Thus, the cost function decreases monotonically at each iteration.
Global Optimality: After solving each subproblem $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$, we compute the reduced costs of all variables. If all reduced costs satisfy $r_{ij}\geq 0$, the algorithm terminates; otherwise, the next iteration begins. By the properties of the reduced cost, this condition guarantees that the final solution is globally optimal. ∎
Proof of Part 2.
The total number of possible combinations of the basis variables among $N$ variables is at most $\binom{N}{M-1}$. Based on the selection rule, in each iteration, the set $\mathcal{I}_{k}$ always contains at least one variable with a negative reduced cost. Denote the variable with the smallest reduced cost as $x_{l}$. When solving the subproblem using the NS method, $x_{l}$ will be added to the basis variables, while one of the current basis variables $\mathcal{B}_{k}$will leave the basis. This implies that each iteration examines at least one new combination of the basis variables. Consequently, after at most $\binom{N}{M-1}$ iterations, all possible combinations of the basis variables will have been examined. Therefore, the algorithm is guaranteed to terminate within a finite number of iterations. ∎
@@ 4.3 Grouped Selection Block Coordinate Descent Network Simplex method
Since an optimal solution to the OT problem is typically sparse, most variables take value zero and do not belong to the optimal basis. Consequently, selecting $\mathcal{I}_{k}$ uniformly at random often results in the inclusion of irrelevant variables, slowing convergence.
To address this issue, we propose a grouped block coordinate selection strategy that partitions the variable index set $\{1,\dots,N\}$ into two disjoint subsets:
  • $\mathcal{R}$: indices of variables deemed likely to enter the optimal basis $\mathcal{B}^{*}$,
  • $\mathcal{S}$: indices of variables with low likelihood of being part of the optimal basis $\mathcal{B}^{*}$.
@@
Find an initial feasible solution $X^{0}\in\mathbb{R}^{n\times m}$. Set the initial basis variables as $\mathcal{J}_{0}$. Choose parameters $s,t\in(0,1)$ such that $s<t$. Initialize $\mathcal{S}=\emptyset$ and =$\mathcal{R}=\{1,2,\ldots,N\}$. Shuffle $\mathcal{R}$ randomly. Define the threshold sequence $\{e_{k}\}$.
@@
     Step 2: Move the non-selected variables with positive reduced costs to $\mathcal{S}$, and return variables with non-positive reduced costs to $\mathcal{R}$.
@@
     Step 6: Optimality check: If $\mathcal{R}=\emptyset$, compute reduced costs for all variables. If $r_{ij}\geq 0$ for all $i,j$, terminate; otherwise, return to Step 1.
@@
Let $\tilde{k}$ denote the total number of outer iterations, and let $l_{k}$ represent the number of NS iterations required to solve the subproblem at iteration $k$. The time complexity of the GS-BCDNS method is given by:
@@
Compared to the Sinkhorn algorithm, which requires $O(N)$ space, the GS-BCDNS method is more memory-efficient. However, relative to existing BCD methods, the memory usage depends on the size of the selected block $\mathcal{H}_{k}$ [xie2023randomized].
@@ 5.   Numerical Experiments
In this section, we perform numerical experiments on various instances of OT problems1.11footnotetext: 1 Numerical experiments were conducted on a computer with an Intel Core i7-12700H CPU (2.70 GHz) and 16 GB of RAM. The implementation was performed using Python 3.8, and the code has been made publicly available on GitHub at https://github.com/842422492/BCDNS.git. We first solve randomly generated OT problems using the GS-BCDNS method to evaluate the impact of parameters $s$ and $t$ on algorithmic efficiency. Subsequently, we compute the Wasserstein distance between uniform and normal distributions using different algorithms to compare their performance.
5.1 Experiment Settings
For simplicity, we assume that the numbers of supply and demand nodes are equal, i.e., $n=m$.
Test Problem 1: Ten randomly generated discrete OT problems using randomly sampled numbers. $n=1000$.
Test Problem 2: The Wasserstein distance between a one-dimensional uniform distribution on $[-1,1]$ and a standard normal distribution on $\mathbb{R}$. $n=1000$.
Test Problem 3: The Wasserstein distance between a two-dimensional uniform distribution on $[-1,1]^{2}$ and a standard normal distribution on $\mathbb{R}^{2}$. $n=1000$.
Test Problem 4: The Wasserstein distance between a three-dimensional uniform distribution on $[-1,1]^{3}$ and a standard normal distribution on $\mathbb{R}^{3}$. $n=1000$.
Test Problem 5: The Wasserstein distance between a one-dimensional uniform distribution on $[-1,1]$ and a standard normal distribution for different problem sizes. $n$ starts from 50 and increases by 50 up to 300.
The discrete OT problems in Test Problem 1 were generated as follows: two sets of $n=1000$ random numbers were sampled from a uniform distribution on $[1,1000]$ to represent the supply and demand distributions $P$ and $Q$, respectively. Then adjuste the two sets to ensure that their sums were equal. Additionally, a cost matrix of size $1000\times 1000$ was generated from the same uniform distribution.
The Wasserstein distances in Test Problem 2, 3, and 4 were estimated as follows: $n$ samples were randomly drawn from the uniform and normal distributions, denoted as $y$ and $\tilde{y}$, respectively. For each pair, i.i.d. samples $\{\tilde{y}_{i}\}$ and $\{y_{j}\}$, $1\leq i,j\leq 1000$, were used to approximate the two continuous probability measures. The cost matrix was defined as: $c_{i,j}=\text{dist}(\tilde{y}_{i},y_{j})^{2},$ and the probability mass vectors were set as $\mathbf{p}=\mathbf{q}=(1/n,\dots,1/n)^{T}\in\mathbb{R}^{n}$.
Figure 1: The point distributions in Test Problem 2, 3 and 4
Algorithm Settings:
  • Sinkhorn (Stabilized) Algorithm: As used in [xie2023randomized], we follow the method proposed in [cuturi2013sinkhorn] for computing the Wasserstein distance. Let $\gamma$ be the entropy regularization coefficient, which is set to $\gamma=1/(4\log n)$ as suggested in [dvurechensky2018computational]. We consider the parameter settings $\varepsilon=10^{-4},10^{-3},0.01,0.1$. Sinkhorn iterations are projected onto the feasible region using a rounding procedure, specifically Algorithm 2 in [altschuler2017near]. To mitigate numerical instability, all updates are performed in log space using the LogSumExp function. The algorithm terminates when either the change in the objective function value falls below $\delta=10^{-6}$ or the runtime exceeds 2000 seconds.
  • GS-BCDNS: When solving Test Problem 1, the parameter $s$ was varied from $1/n$ to 0.9, while the parameter $t$ was varied from $2/n$ to 1. For Test Problems 2, 3, 4 and 5, the values of $s$ and $t$ were fixed at $5/n$ and $10/n$, respectively. For each Test Problem, set $\{e_{k}\}$ to positive infinity and calculate the initial basic feasible solution using the Northwest Corner Rule.

### v3

@@ 4.  Block Coordinate Network Simplex Methods with Basis Variable Succession
  • $\mathbf{x}^{*}$:the optimal solution to the OT problem (2),
@@ 4.1 Basis Variables Succession
The following theorem clarifies when solving a subproblem with the NS method restricted to $\mathcal{H}_{k}$ can yield a globally optimal solution of the full OT problem, and when this is impossible.
@@
Apply the NS method to the subproblem restricted to $\mathcal{H}_{k}$, and set
@@
    If $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$ and $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, then $\mathbf{x}^{k+1}$ is a globally optimal solution of the full OT problem.
@@
    Assume that the global optimal solution $\mathbf{x}^{\ast}$ is unique. If $\mathcal{B}_{k}\not\subseteq\mathcal{H}_{k}$, then even when $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the iterate $\mathbf{x}^{k+1}$ cannot be globally optimal.
@@
(i) Because $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$, the current basic feasible solution $\mathbf{x}^{k}$ is feasible for the restricted subproblem, and thus the NS method can be initialized from $\mathbf{x}^{k}$ without losing feasibility. Let $\mathbf{x}^{k+1}$ be an optimal solution of the restricted subproblem. Then $\mathbf{x}^{k+1}$ is feasible for the full OT problem (it satisfies $A\mathbf{x}^{k+1}=b$ and $\mathbf{x}^{k+1}\geq 0$), hence
$c^{\top}\mathbf{x}^{k+1}\ \geq\ c^{\top}\mathbf{x}^{\ast}.$
On the other hand, since $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the basic feasible solution $\mathbf{x}^{\ast}$ (with nonbasic variables equal to zero) is feasible for the restricted subproblem as well. By optimality of $\mathbf{x}^{k+1}$ for the restricted subproblem, we have
$c^{\top}\mathbf{x}^{k+1}\ \leq\ c^{\top}\mathbf{x}^{\ast}.$
Combining the two inequalities yields $c^{\top}\mathbf{x}^{k+1}=c^{\top}\mathbf{x}^{\ast}$, so $\mathbf{x}^{k+1}$ attains the global optimal value of the full OT problem. Therefore $\mathbf{x}^{k+1}$ is globally optimal.
(ii) Assume $\mathbf{x}^{\ast}$ is the unique globally optimal solution. Let $j\notin\mathcal{H}_{k}$ with $x^{k}_{j}>0$. By the definition of the restricted subproblem, the $j$-th component cannot change:
$x^{k+1}_{j}\;=\;x^{k}_{j}+d^{k}_{j}\;=\;x^{k}_{j}\;>\;0.$
Now suppose (for contradiction) that $\mathbf{x}^{k+1}$ is globally optimal. By uniqueness of the global optimizer, this would imply $\mathbf{x}^{k+1}=\mathbf{x}^{\ast}$, and in particular $x^{\ast}_{j}=x^{k+1}_{j}>0$.
However, under the assumption $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, we have $j\notin\mathcal{B}^{\ast}$, and since $\mathcal{B}^{\ast}$ is a basis of a basic feasible optimal solution, all variables outside $\mathcal{B}^{\ast}$ are nonbasic and therefore equal to zero. Thus $x^{\ast}_{j}=0$, contradicting $x^{\ast}_{j}>0$. Consequently, $\mathbf{x}^{k+1}\neq\mathbf{x}^{\ast}$, and $\mathbf{x}^{k+1}$ cannot be globally optimal. ∎
The above result motivates a deterministic working set construction rule, which we call basis-variable succession. Specifically, at each iteration $k$, we choose the block $\mathcal{H}_{k}$ so that it contains the current basis, i.e., $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$. This choice contrasts with existing BCD-type approaches [18], where the working set is generated without explicitly preserving the current basis.
From a computational perspective, let $L_{k}$ denote the number of NS pivots required to solve the restricted subproblem over $\mathcal{H}_{k}$, and let $L$ denote the pivot count when applying the NS method to the full problem. Then the dominant cost of the restricted solve scales as $O(L_{k}\,|\mathcal{H}_{k}|)$, which can be substantially smaller than the $O(LN)$ cost incurred by scanning all $N$ variables in each pivot of the full problem.
This strategy has two immediate practical benefits. First, since the restricted subproblem retains all equality constraints and contains at least M1M-1 variables, it admits a feasible spanning tree basis and hence remains well-posed throughout the iterations. Second, reusing $\mathcal{B}_{k}$ as a warm start provides a readily available initial basis for the next subproblem, avoiding the overhead of repeatedly constructing an initial feasible basis from scratch.
@@ 4.2 A Prototype of Block Coordinate Descent Network Simplex Methods and its Finite Termination
Following the strategy of section 4.1, choose a block
@@
Let $q(\mathbf{x}^{k};\mathcal{H}_{k})$ denote the optimal value of the restricted subproblem $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$. The next theorem establishes finite termination and global optimality of the resulting sequence $\{\mathbf{x}^{k}\}$.
@@
Consider the prototype BCDNS method with the block selection rule (7). The sequence $\{q(\mathbf{x}^{k};\mathcal{H}_{k})\}_{k\geq 0}$ is non-increasing monotone. Moreover, if the NS subproblem solver is equipped with Bland’s anti-cycling pivot rule [4], then the method after finitely many degenerate pivots, either a strictly improving pivot occurs or the method terminates. In particular, the algorithm terminates in finitely many iterations and returns a globally optimal OT solution.
Proof.
Fix an outer iteration $k$ and consider the working set $\mathcal{H}_{k}=\mathcal{B}_{k}\cup\mathcal{I}_{k}$. We solve the restricted subproblem $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$ by the NS method started on the current basis $\mathcal{B}_{k}$.
Monotonicity (non-increase). A basic property of the NS method is that each pivot maintains feasibility and does not increase the objective value; hence the optimal value of the restricted subproblem satisfies
$q(\mathbf{x}^{k};\mathcal{H}_{k})\;\leq\;q(\mathbf{x}^{k};\mathcal{B}_{k}).$ (8)
Since $\mathcal{B}_{k}$ is the optimal basis obtained from the previous restricted subproblem, the start objective equals the previous subproblem optimum:
$q(\mathbf{x}^{k};\mathcal{B}_{k})\;=\;q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$ (9)
Combining (8)–(9) yields
$q(\mathbf{x}^{k};\mathcal{H}_{k})\;\leq\;q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}),$
so $\{q(\mathbf{x}^{k};\mathcal{H}_{k})\}$ is non-increasing monotone.
Finite termination and eventual strict decrease under Bland’s rule. Degenerate pivots may yield zero objective improvement, so strict decrease cannot be guaranteed at every pivot. However, when Bland’s anti-cycling pivot rule is used in the NS solver [4], cycling is prevented and hence a basis cannot repeat. Since the number of feasible bases is finite, each restricted NS solve terminates in finitely many pivots; moreover, it cannot perform infinitely many zero-improvement pivots, i.e., after finitely many degenerate pivots it either makes a strictly improving pivot or reaches restricted optimality.
At the outer iteration, if the negative reduced cost set $\mathcal{N}_{k}$ is empty, then all reduced costs are nonnegative and $\mathbf{x}^{k}$ is globally optimal, so the method terminates. Otherwise, $\mathcal{N}_{k}\neq\varnothing$ and the block selection rule (7) guarantees that the restricted subproblem contains at least one negative reduced cost index; therefore the NS solver of the restricted subproblem at iteration $k$ is well defined and terminates after finitely many pivots. Consequently, the overall method terminates in finitely many outer iterations.
4.3 Random Selection Block Coordinate Descent Network Simplex Method
A natural way to instantiate the prototype in section 4.2 is to construct the working set by random sampling. we construct at iteration $k$ a working set $\mathcal{H}_{k}\supseteq\mathcal{B}_{k}$ by augmenting the current basis $\mathcal{B}_{k}$ with a randomly sampled subset $\mathcal{I}_{k}$. The key requirement inherited from (7) is that the block must contain at least one index with negative reduced cost whenever the current iterate is not yet globally optimal.
Fix a sampling ratio $s\in(0,1)$ and set $b:=\lvert\mathcal{B}_{k}\rvert=M-1$. At iteration $k$, we draw $\lvert\mathcal{I}_{k}\rvert=\lceil sN\rceil$ indices uniformly at random from the complement of the current basis, $\{1,\dots,N\}\setminus\mathcal{B}_{k}$. We compute the reduced costs of the sampled indices and enforce the descent condition
$\mathcal{I}_{k}\cap\mathcal{N}_{k}\neq\varnothing,\qquad\mathcal{N}_{k}:=\{\ell:\bar{c}^{k}_{\ell}<0\}.$
Then we form
$\mathcal{H}_{k}:=\mathcal{B}_{k}\cup\mathcal{I}_{k}\quad\text{and solve}\quad\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})\ \text{by the NS method.}$
The algorithm is presented below.
Algorithm 1 Random Selection Block Coordinate Descent Network Simplex Method (RS-BCDNS)
Initialization: Find an initial feasible solution $\mathbf{x}^{0}$ and an initial basis $\mathcal{B}_{0}$ Choose a sampling ratio $s\in(0,1)$.
for $k=0,1,2,\dots$ do
  Compute reduced costs for all indices. If $\bar{c}^{k}_{\ell}\geq 0$ for all $\ell=1,\dots,N$ terminate
  repeat
   Sample $\mathcal{I}_{k}\subseteq\{1,\dots,N\}\setminus\mathcal{B}_{k}$ uniformly at random with $\lvert\mathcal{I}_{k}\rvert=\lceil sN\rceil$.
   Compute $\bar{c}^{k}_{\ell}$ for all $\ell\in\mathcal{I}_{k}$.
  until $\min_{\ell\in\mathcal{I}_{k}}\bar{c}^{k}_{\ell}<0$
  Set $\mathcal{H}_{k}\leftarrow\mathcal{B}_{k}\cup\mathcal{I}_{k}$.
  Solve $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$ by the NS method from $\mathcal{B}_{k}$, obtaining $d^{k}$ and the updated basis $\mathcal{B}_{k+1}$.
  Update $\mathbf{x}^{k+1}\leftarrow\mathbf{x}^{k}+d^{k}$.
end for
@@ 4.4 Grouped Selection Block Coordinate Descent Network Simplex method
In section 4.3, we introduced a random selection variant as a natural baseline. Although this random augmentation is simple and preserves the basis-succession mechanism, it can be highly inefficient for large-scale OT instances. Indeed, optimal transport plans are typically sparse, and only a small fraction of variables eventually belong to an optimal basis. Consequently, sampling $\mathcal{I}_{k}$ uniformly at random tends to include many variables that are unlikely to become basis, so their reduced costs are evaluated repeatedly and they may participate in subproblems without contributing to meaningful progress.
To this end, we propose a grouped block-coordinate selection strategy that partitions the index set $\{1,\dots,N\}$ into two disjoint subsets:
  • $\mathcal{R}$: indices that are deemed likely to enter an optimal basis $\mathcal{B}^{\ast}$,
  • $\mathcal{S}$: indices with low likelihood of being part of $\mathcal{B}^{\ast}$.
@@
Find an initial feasible solution $X^{0}\in\mathbb{R}^{n\times m}$. Set the initial basis variables as $\mathcal{J}_{0}$. Choose parameters $s,t\in(0,1)$ such that $s<t$. Initialize $\mathcal{S}=\emptyset$ and $\mathcal{R}=\{1,2,\ldots,N\}$. Shuffle $\mathcal{R}$ randomly. Define the threshold sequence $\{e_{k}\}$.
@@
  Step 2: Move the non-selected variables to $\mathcal{S}$.
@@
  Step 6: Optimality check: If $\mathcal{R}=\emptyset$, compute reduced costs for all variables.
  if $r_{ij}\geq 0$ for all $i,j$ then
   terminate.
  else
   Update $\mathcal{R}\leftarrow\{(i,j):r_{ij}<0\}$.
   return to Step 1.
  end if
@@
Let $\tilde{k}$ denote the total number of iterations, and let $l_{k}$ represent the number of NS iterations required to solve the subproblem at iteration $k$. The time complexity of the GS-BCDNS method is given by:
@@
Compared to the Sinkhorn algorithm, which requires $O(N)$ space, the GS-BCDNS method is more memory-efficient.
@@ 5.  Numerical Experiments
In this section, we report numerical experiments designed to (i) study the effect of the block-size parameters $(s,t)$ in the proposed BCDNS framework, (ii) compare the proposed methods with representative baselines on the Wasserstein distance computations, and (iii) evaluate the performance and scalability of the proposed methods on large-scale OT instances.11footnotetext: Numerical experiments were conducted on a computer with an Intel Core i7-12700H CPU (2.70 GHz) and 16 GB of RAM. The implementation was performed in Python 3.8. The code and the scripts for reproducing all figures are publicly available at https://github.com/842422492/BCDNS.git.
5.1 Experimental Setup
This section describes the common procedure used to generate OT test instances and the implementation details shared by all methods in Section 5.2–Section 5.4. Unless otherwise stated, we consider balanced OT with equal numbers of supply and demand nodes, i.e., $n=m$.
Test instance generation and reproducibility.
All Wasserstein test instances are constructed by Monte Carlo discretization of continuous probability measures. For a prescribed sample size $n$, we draw i.i.d. samples $\{u_{i}\}_{i=1}^{n}$ from a source distribution $\mu$ and $\{v_{j}\}_{j=1}^{n}$ from a target distribution $\nu$. We define the quadratic transport cost
$c_{ij}\;=\;\|u_{i}-v_{j}\|^{2},\qquad 1\leq i,j\leq n,$ (10)
and set the discrete marginals to be uniform,
$p=q=\Bigl(\tfrac{1}{n},\dots,\tfrac{1}{n}\Bigr)^{\top}\in\mathbb{R}^{n}.$ (11)
The resulting linear program is then solved by the competing methods. To ensure reproducibility, we save and version-control all sampled point sets $\{u_{i}\}$ and $\{v_{j}\}$ for every reported instance; hence the corresponding cost matrix $C$ in (10) can be reconstructed exactly.
Algorithmic settings.
  • Sinkhorn (stabilized). Following [6] and the stabilized implementation used in [18], we solve the entropically regularized OT problem with regularization coefficient $\gamma=\varepsilon/(4\log n)$ as suggested in [7]. We test $\varepsilon\in\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. Each Sinkhorn iterate is rounded to feasibility using the projection/rounding procedure (Algorithm 2 in [2]). To improve numerical stability, all updates are carried out in log-domain via the LogSumExp operator. The algorithm terminates when the absolute change in objective value falls below $\delta=10^{-6}$ or when the runtime exceeds $2000$ seconds.
  • GS-BCDNS.
    Our NS method is implemented by following the design of the network simplex solver in the Python NetworkX package. In the experiments, we multiply the cost matrix by a positive scaling factor and round to integers to obtain an equivalent integer-cost instance for the NS pivots; this scaling does not change the set of optimal transport plans, since it preserves the ordering of feasible solutions by objective value. In Section 5.2, we first perform a search on $(s,t)$ to identify an efficient regime for subproblem sizes. Based on the empirical findings in Section 5.2, we fix $(s,t)$ for all subsequent comparisons. Unless otherwise stated, we use
    $s=\tfrac{2}{n},\qquad t=\tfrac{20}{n}.$
  • RS-BCDNS. The method is initialized by the Northwest Corner method and set the block size to 256.